Question 18

A deck of cards contains 12 green 12 red and 12 white cards. What is the probability (stated as a percentage) of randomly selecting a green card from this group?

cannot be determined
93.75%
33.33%
66.67%
Answer #1

P(Randomly selecting a green card)

= 12 / (12 + 12 + 12)

= 12/36

= 1/3

= 0.3333

= 33.33%

Option C is correct.